Evaluate the iterated Integral  in terms of e. 


1
Expert's answer
2021-05-24T16:08:30-0400

For example:

"\displaystyle \iint_Re^{xy}dA,\ R=[2,4]\times [1,2]"


Then:

"\displaystyle \iint_Re^{xy}dA=\displaystyle\int^4_2\int^2_1e^{xy}dydx=\displaystyle\int^4_2(e^{xy}|^2_1)dx="

"\displaystyle\int^4_2(e^{2x}-e^x)dx=(e^{2x}/2-e^x)|^4_2=e^8/2-e^4-e^4/2+e^2=e^8/2-3e^4/2+e^2"